A Complete Bibliography of ACM Transactions on Internet of Things (TIOT)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

21 December 2023 Version 1.13

Title word cross-reference

3 [ASD^{+}20]. 5 [AEMK^{+}21].

-facilitated [ZUH^{+}22]. -Modeling [AEMK^{+}21].

2 [TDD^{+}22].

6G [PF23]. 6TiSCH [KK21, KK22, RVDA20].

Access [SBC^{+}22]. Accurate [MKC^{+}23]. Achieving [LXL^{+}23]. ACM [DP20]. Activities [JPGM20, SLLC21]. Activity [LMX^{+}20, MAM^{+}21, YNS^{+}21]. Adaptation [HH22]. Adaptive [NLQ22, KKR^{+}22]. Adversarial [YNS^{+}21]. Aerial [BMPM20]. Affect [AENM^{+}22]. Agile [MBHF23]. agnostic [ZZL^{+}23]. Agriculture [BMPM20]. AI [MAD^{+}22, DSS^{+}20]. Air [LHJW23, LZW^{+}23]. airBP [LZW^{+}23]. Algorithms [DGL^{+}21, DSS^{+}20]. All-wireless [MCRP20]. Allocation [NG21, XDP^{+}22]. Analysis [ATB^{+}22, BRL22]. Analytics [BMPP21]. Annotated [GTG^{+}21]. Anomaly [NLQ22, SB21]. Anonymizing [HAK22]. Application [PDM^{+}20]. Applications [GMT23, KKR^{+}22, KLMS22, XDP^{+}22]. Approach [GGBD20, HAK22, KWS22, NLQ22, PF23, SB21]. Architecture [ND20, PMZ23]. Arrival [MKC^{+}23]. Assignment [BPR^{+}23]. Assisted [AAA^{+}21]. ASTRO [PSLK21]. Attack [CRCR21]. Attestable [PSW^{+}22]. Attestation [AC20]. Audio [TF21]. Augmented [ZUH^{+}22]. Authenticating [JLZ^{+}20]. Authentication [ARA23, KKBL20, SLLC21]. Authorization [KKBL20]. Automated [KKR^{+}22]. Automatic [PDM^{+}20]. Automation [CKCZY22]. Autonomic [LKWM21]. Autonomous [BMPM20]. Aware [SAGT20, SBC^{+}22, XDP^{+}22, BJM21, MGSD22, NG21, PF23, TFPNT23]. Awareness [AWRP23, WLL^{+}21].

Backscatter [KHG^{+}20]. Bandit [NLQ22]. Based [ATB^{+}22, AAA^{+}21, BMPM20, BRL22, BBS^{+}21, CRCR21, FAlS^{+}22, ZYJ^{+}23, ARA23, BA20, GGBD20, HST^{+}20, LMX^{+}20, NG21, WLL^{+}21, ZZL^{+}23, KKY22]. Battery [PCMS21]. Battery-free [PCMS21]. Behaviour [NG21]. Behaviour-aware [NG21]. Bend [ATB^{+}22]. Best [MBHF23]. between [jWCS21]. BLE [MCRP20]. Bleeding [LLL^{+}22]. Blockchain [ARA23]. Blood [LZW^{+}23]. BlueFlood [AEMK^{+}21]. Bluetooth [AEMK^{+}21, SBR20]. Bonseyes [DSS^{+}20]. Bringing [DSS^{+}20]. Buildings [CKCZY22, PCMS21, SB21]. bus [MKC^{+}23]. Buses [KKY22].

Cache [HS21]. Camera [HNB^{+}20]. Cameras [KQTC20]. Capacitor [LMX^{+}20]. Capacitor-based [LMX^{+}20]. Capsule [LLL^{+}22]. Capture [PSW^{+}22]. carried [BPR^{+}23]. Case [AAA^{+}21, HGG^{+}23]. CH [LXL^{+}23]. CH-MAC [LXL^{+}23]. Channel [FAlS^{+}22, PF23]. Channel-aware [PF23]. Channels [JLZ^{+}20]. Characterization [MKC^{+}23]. Charging [EAM^{+}21, YXH^{+}22]. Chart [LWLJ22]. Chief [Pic23]. Chimera [ADY^{+}21]. Cities [ZCGL22]. City [HGG^{+}23, MKC^{+}23, MAD^{+}22]. City-bus [MKC^{+}23]. Class [BRL22]. Classification [ACS20, BRL22, GGBD20, HST^{+}20, KWS22]. Closed [ZGX^{+}23]. Closed-loop [ZGX^{+}23]. Cloud [TPFF20]. Cloud-smart [TPFF20]. Clouds [ASD^{+}20, DLR^{+}23, HS21, MAD^{+}22]. CoCo [jWCS21]. Coding [LXL^{+}23]. Coil [YXH^{+}22]. Collaborative [JTLL21, XDP^{+}22]. Collection [TPFF20]. Communication [GMHT22, LXL^{+}23, PCMS21, SAGT20, TDB^{+}23]. Communications [MGSD22]. Comparison [GMT23]. Components [GGBD20]. Comprehensive [GMT23]. Compressed [ZCC^{+}21]. Compromised [BEM22]. Computer [BDJ^{+}22]. Computing [DLR^{+}23, GGT^{+}23, KKBL20, MVC22, MZC^{+}22, NG21, NPS^{+}22, NLQ22]. Concurrent [AEMK^{+}21]. Connected [HNB^{+}20, AENM^{+}22]. Conquer [GGBD20]. Considerations [KSS21]. Constrained [KWS22, ACS20]. Consumer [AENM^{+}22]. Consumption [KSBD20]. Context [WLL^{+}21]. Contextual [MKC^{+}23, NLQ22]. Contextual-Bandit [NLQ22]. Contextualizing [EAM^{+}21]. Continual [YNS^{+}21]. Continuous [BMPP21, KHG^{+}20]. Control [AWRP23, AAA^{+}21, KK21, LWLJ22, SBC^{+}22]. Controlling [SBC^{+}22]. Cooperative [GGT^{+}23]. Copresence [FAlS^{+}22]. Correct [KLMS22]. Correlations [jWCS21]. Cost [DLR^{+}23, BA20]. Could [LKR^{+}23]. Coverage [HNB^{+}20]. CPS [PDM^{+}20]. criteria [NG21]. Crop [BA20]. Crowd [ASD^{+}20]. Crowd-sourced [ASD^{+}20]. Crowdsensing [ZCGL22]. Cryptocurrency [ZGX^{+}23]. Cryptosystem [ARA23]. Cryptosystem-based [ARA23]. Cube [ARA23]. Cycle [BJM21].

D [ASD^{+}20]. Daily [SLLC21]. Data [BRL22, BBS^{+}21, BMPP21, DLR^{+}23, HAK22, HGG^{+}23, PSW^{+}22, WLL^{+}21, jWCS21, WLS23, ZGX^{+}23, DSS^{+}20]. Deep [MAM^{+}21, SLLC21]. Deeply [ZCC^{+}21]. Demarcation [BMPM20]. Deployment [HNB^{+}20, KKR^{+}22, KLMS22, DSS^{+}20]. Deployments [NCV^{+}22]. Design [AAN^{+}23, EFO^{+}20, KLMS22]. Detection [CRCR21, FAlS^{+}22, HST^{+}20, JCY^{+}23, LLL^{+}22, NLQ22, SB21, ZYJ^{+}23]. Development [LKR^{+}23, PDM^{+}20]. Device [DGL^{+}21, SBC^{+}22, WEOR23, WLL^{+}21, YXH^{+}22]. Device-free [WLL^{+}21]. Devices [ACS20, AENM^{+}22, BPR^{+}23, JLZ^{+}20, TDD^{+}22, ZUH^{+}22, ZCC^{+}21]. Different [GGBD20]. Digital [BDJ^{+}22, PMZ23]. Distributed [RVDA20, TDD^{+}22, DLT^{+}23]. Diverse [NCV^{+}22]. Divide [GGBD20]. Divide-and-Conquer-based [GGBD20]. Domain [HH22]. Driven [ARA23, KSBD20, KKR^{+}22, PCMS21]. Driving [JCY^{+}23, KSS21]. Drone [PSLK21]. DSME [GMT23]. Duty [BJM21]. Dynamic [KKY22, NG21, SBR20, TFPNT23]. Dynamics [BRL22, KSS21].

Early [AAA^{+}21, GGBD20]. Ease [ZUH^{+}22]. Ecosystems [PF23]. Edge [AAA^{+}21, BMPP21, DLT^{+}23, GGT^{+}23, HAK22, HS21, KKBL20, MVC22, MAD^{+}22, NLQ22, PF23, PMZ23, TFPNT23]. Edge-Assisted [AAA^{+}21]. Editor [Pic23]. Editor-in-Chief [Pic23]. Editorial [DP20]. Efficient [ACS20, BBS^{+}21, TDB^{+}23]. Elastic [TFPNT23]. Electric [EAM^{+}21, KSS21]. Elk [TF21]. Enabled [KK22, MAD^{+}22, SLLC21]. Enabling [HS21]. End [DSS^{+}20]. End-to-end [DSS^{+}20]. Endoscopy [LLL^{+}22]. Energy [KSBD20, SBR20, TDB^{+}23, XDP^{+}22]. Energy-Aware [XDP^{+}22]. Energy-Efficient [TDB^{+}23]. Enhancing [AWRP23, MBHF23]. Entities [EAM^{+}21]. Environment [ARA23, PCMS21, SBC^{+}22]. Environment-driven [PCMS21].

Environmental [BK21]. Environments [BMPP21, KWS22, PSW^{+}22, SBR20]. Evaluation [AEMK^{+}21, EFO^{+}20, GMHT22, PMZ23, RVDA20]. Event [BEM22, BBS^{+}21]. Exchange [BBS^{+}21]. Exercises [HST^{+}20]. Exploiting [MKC^{+}23]. Expression [ZCC^{+}21]. Expressive [KLMS22]. Extending [ND20].

Facial [ZCC^{+}21]. facilitated [ZUH^{+}22]. Fast [ZCC^{+}21]. Faster [KK21]. Fatigue [JCY^{+}23]. Fault [NCV^{+}22]. Feature [ZYJ^{+}23]. Features [LLL^{+}22]. Federated [SB21, WEOR23]. Fine [JCY^{+}23, PDM^{+}20]. Fine-grained [JCY^{+}23, PDM^{+}20]. Fingerprinting [SDS^{+}22, WEOR23]. FL [PF23]. FL4IoT [WEOR23]. Fleet [KKY22]. Flexible [NCV^{+}22, PMZ23]. FlockLab [TDD^{+}22]. Floorplans [GTG^{+}21]. Fog [DLR^{+}23, LWLJ22, MZC^{+}22, NG21]. Formation [KK21, KK22]. Framework [BPR^{+}23, GGT^{+}23, MCRP20, WLL^{+}21, ZGX^{+}23]. free [PCMS21, WLL^{+}21]. Frequency [ZYJ^{+}23]. Fusing [WLS23]. Fusion [LLL^{+}22].

Gait [ATB^{+}22, LWL^{+}20]. General [WLL^{+}21]. Generation [ARA23]. Generative [YNS^{+}21]. Geo [DLT^{+}23]. Geo-distributed [DLT^{+}23]. Geolocation [LXML21]. Goalkeeper [HST^{+}20]. grained [JCY^{+}23, PDM^{+}20]. Grammar [ASD^{+}20]. Grammar-supported [ASD^{+}20]. Green [CKCZY22]. Grid [EAM^{+}21, PSLK21]. Grief [LKR^{+}23]. Group [GMHT22]. Guitar [TPFF20].

Handling [NCV^{+}22]. Handwritten [LHJW23]. Harvesting [SAGT20]. Harvesting-Aware [SAGT20]. Haunted [BEM22]. Health [EFO^{+}20]. Healthcare [AAA^{+}21]. Heterogeneous [XDP^{+}22]. Hierarchical [NLQ22]. Home [BEM22, CKCZY22, JLZ^{+}20, SBC^{+}22]. Homes [KSBD20]. hop [AEMK^{+}21]. Hopping [LXL^{+}23]. House [BEM22]. Human [BPR^{+}23]. Human-carried [BPR^{+}23]. Hybrid [ZGX^{+}23, ZCGL22]. HyEdge [GGT^{+}23].

IDeA [BMPM20]. Identification [BPR^{+}23, LWL^{+}20, SDS^{+}22, WEOR23]. IDIoT [BPR^{+}23]. IEEE [GMT23]. Images [LLL^{+}22]. Implementation [PMZ23]. Improving [SBR20]. In-Air [LHJW23]. Inaugural [DP20]. Independent [JTLL21, MAM^{+}21]. Indoor [ASD^{+}20, GTG^{+}21, JPGM20]. Industrial [GMT23]. Inertial [ATB^{+}22]. Inference [ASD^{+}20]. Information [FAlS^{+}22]. Infrastructure [SAGT20]. Infrastructure-Less [SAGT20]. Infrastructures [MZC^{+}22]. Inhomogeneous [TDB^{+}23]. Insider [CRCR21]. inspired [ZJZ^{+}23]. Instrument [TPFF20]. Integrating [ZCGL22]. integration [DSS^{+}20]. Intelligent [AC20, PF23]. Interactions [KHG^{+}20, TPFF20]. Interactive [AWRP23]. Interface [HGG^{+}23]. Internet [DP20, AWRP23, ARP21, AC20, AAA^{+}21, AENM^{+}22, ADY^{+}21, BRL22, BJM21, CRCR21, GMHT22, HGG^{+}23, JTLL21, KKBL20, KKR^{+}22, MGSD22, NLQ22, TF21, ZUH^{+}22]. Internet-connected [AENM^{+}22]. Interoperability [ND20]. Interpersonal [AENM^{+}22]. Introduction [MHL^{+}23]. Intrusive [TDD^{+}22]. IoT [KKBL20, ZUH^{+}22, ACS20, AENM^{+}22, ARA23, BMPM20, BBS^{+}21, CKCZY22, EFO^{+}20, GMT23, GGBD20, JPGM20, KK22, KQTC20, KKY22, LKWM21, LKR^{+}23, MHL^{+}23, MVC22, MBHF23, NCV^{+}22, ND20, PSW^{+}22, Pic23, SDS^{+}22, TDD^{+}22, TDB^{+}23, WEOR23, XDP^{+}22, ZGX^{+}23, ZCGL22]. IoT-Based [KKY22, BMPM20]. IoT-Sensing [ZCGL22]. IoTRepair [NCV^{+}22]. IoTs [LMX^{+}20]. IP [LXML21]. IPv6 [KK22]. Issue [DP20, MHL^{+}23, Pic23].

Key [ARA23]. Keystroke [ZZL^{+}23]. Kinetic [LMX^{+}20]. Kinetic-powered [LMX^{+}20].

LA-MQTT [MGSD22]. LandmarkMiner [LXML21]. Landmarks [LXML21]. Languages [LKR^{+}23]. Large [TPFF20]. latency [LXL^{+}23]. Learning [CRCR21, DGL^{+}21, HAK22, HH22, JTLL21, MAM^{+}21, SB21, SLLC21, WEOR23]. Less [SAGT20]. level [LXML21]. Leveraging [WWZZ20]. Lights [WWZZ20]. Lightweight [AC20, JTLL21]. Limited [JLZ^{+}20]. Living [BMPP21]. Locating [KQTC20]. Location [BJM21, MAM^{+}21, MKC^{+}23, MGSD22]. Location- [MAM^{+}21]. Location-aware [MGSD22]. loop [ZGX^{+}23]. LoRa [AAN^{+}23]. Low [ADY^{+}21, BA20, DLR^{+}23, LXL^{+}23, SBR20]. Low-Cost [DLR^{+}23, BA20]. Low-latency [LXL^{+}23]. Low-power [ADY^{+}21]. LPWAN [LXL^{+}23]. LSR [TDB^{+}23]. LSTM [ZCC^{+}21].

MAC [AAN^{+}23, LXL^{+}23]. Machine [CRCR21, DGL^{+}21, PF23]. MAIDE [ZUH^{+}22]. Maintenance [KKY22]. MakeSense [JPGM20]. Management [AWRP23, LKWM21, MZC^{+}22, RVDA20]. Managing [XDP^{+}22]. Manufacturing [LWLJ22]. MapSense [ASD^{+}20]. Maritime [MVC22]. Marketplaces [HGG^{+}23]. Me [AENM^{+}22]. Measurement [KHG^{+}20]. Mesh [MCRP20]. Method [LXML21]. Metrics [LWLJ22]. MFD [ZYJ^{+}23]. MGait [ATB^{+}22]. Micro [DLR^{+}23]. Micro-Clouds [DLR^{+}23]. Middleware [BBS^{+}21, MZC^{+}22]. Millimeter [LHJW23, LZW^{+}23].

Millimeter-Wave [LHJW23, LZW^{+}23]. Mining [LXML21]. Missions [PSLK21]. mmDrive [JCY^{+}23]. mmHSV [LHJW23]. mmWave [JCY^{+}23, ZJZ^{+}23]. Mobile [YXH^{+}22, ZUH^{+}22, ZCC^{+}21]. Mobility [jWCS21]. modal [MKC^{+}23]. Mode [RVDA20]. Model [ATB^{+}22, ARA23, KKR^{+}22]. Model-Based [ATB^{+}22]. Model-driven [KKR^{+}22]. Modeling [AEMK^{+}21]. Modification [KKR^{+}22]. Modular [PMZ23]. Modulation [TDB^{+}23]. Monitor [LZW^{+}23]. Monitoring [BA20, LWLJ22, ZJZ^{+}23]. MQTT [MGSD22]. Multi [AEMK^{+}21, LLL^{+}22, MKC^{+}23, NG21, SBC^{+}22, TDB^{+}23, YXH^{+}22, ZYJ^{+}23, SBC^{+}22]. Multi-Coil [YXH^{+}22]. Multi-criteria-based [NG21]. Multi-Device-Aware [SBC^{+}22]. Multi-hop [AEMK^{+}21]. Multi-modal [MKC^{+}23]. Multi-Modulation [TDB^{+}23]. Multi-object [ZYJ^{+}23]. Multi-scale [LLL^{+}22]. Multi-User [SBC^{+}22]. Multimodal [BPR^{+}23]. Multivariate [GGBD20]. Music [TPFF20]. Musical [TPFF20, TF21]. My [SBC^{+}22].

Need [WLL^{+}21]. Network [HNB^{+}20, KK21, KK22, LXML21, ZCC^{+}21]. Networked [PSLK21]. Networks [BJM21, EFO^{+}20, MAM^{+}21, TDB^{+}23, YNS^{+}21]. Neural [MAM^{+}21, ZCC^{+}21]. Next2You [FAlS^{+}22]. No [WLL^{+}21]. Nodes [DLT^{+}23]. Non [TDD^{+}22]. Non-Intrusive [TDD^{+}22]. Nonlinear [HH22]. Notary [PSW^{+}22]. Novel [CRCR21, KWS22].

object [ZYJ^{+}23]. Objects [ASD^{+}20, GTG^{+}21]. Oceans [DLR^{+}23]. Off [PSLK21]. Off-grid [PSLK21]. On-Device [DGL^{+}21]. Onboarding [ZUH^{+}22]. Online [ACS20]. Open [KK22, TF21]. Operating [TF21]. Opportunistic [KK21]. Optimal [SAGT20]. Optimization [KSS21, KSBD20]. OSCORE [GMHT22]. Osmotic [NPS^{+}22].

Packets [KK21]. Passive [WWZZ20]. PassiveVLP [WWZZ20]. Path [BMPM20]. Perceived [KSBD20]. Perceived-Value-driven [KSBD20]. Performance [GMT23, GMHT22]. Person [MAM^{+}21]. Person-Independent [MAM^{+}21]. Perspective [BDJ^{+}22, DGL^{+}21]. Pervasive [DLR^{+}23]. Physical [BEM22, KHG^{+}20]. Physiology [ZJZ^{+}23]. Physiology-inspired [ZJZ^{+}23]. Pi [ZJZ^{+}23]. Pi-ViMo [ZJZ^{+}23]. Pipeline [DSS^{+}20]. Pipeline-Bringing [DSS^{+}20]. Placement [PF23, ZZL^{+}23]. Placement-agnostic [ZZL^{+}23]. Placementwith [TFPNT23]. Planning [BMPM20, CKCZY22]. Platform [ADY^{+}21]. Point [ASD^{+}20]. Positioning [WWZZ20]. Power [EAM^{+}21, ADY^{+}21]. powered [LMX^{+}20]. PPG [AAA^{+}21]. PPG-Based [AAA^{+}21]. Practices [MBHF23]. Pre [WLL^{+}21]. Pre-processing [WLL^{+}21]. Precision [BMPM20]. Prediction [MKC^{+}23]. Predictions [PDM^{+}20]. Predictive [KKY22]. Presence [BEM22]. preserved [MAD^{+}22]. Preserving [KQTC20, ZCGL22, JTLL21]. Pressure [LZW^{+}23]. PrioDeX [BBS^{+}21]. Prioritization [BBS^{+}21]. Privacy [AWRP23, ARP21, BJM21, JTLL21, KQTC20, MAD^{+}22, ZCGL22]. Privacy-aware [BJM21]. Privacy-preserved [MAD^{+}22]. Privacy-Preserving [ZCGL22, JTLL21]. Private [GGT^{+}23]. processing [WLL^{+}21]. Production [BDJ^{+}22]. Profiling [SDS^{+}22]. Project [MBHF23]. Projections [JTLL21]. Proposals [KKR^{+}22]. Protocol [AAN^{+}23]. Provisioning [GGT^{+}23, MZC^{+}22, TFPNT23]. Proximity [TFPNT23]. Proximity-aware [TFPNT23]. Public [GGT^{+}23, KKY22]. Publish [MGSD22]. Publish-subscribe [MGSD22]. Pure [KSS21].

Qi [YXH^{+}22]. Qi-Wireless [YXH^{+}22]. QID [YXH^{+}22]. QQ [LWLJ22]. Qualitative [LWLJ22]. Quantifying [jWCS21]. Quantitative [LWLJ22]. Query [HGG^{+}23]. Querying [TPFF20].

Radar [JCY^{+}23, LHJW23]. Radars [ZJZ^{+}23]. Radio [WLL^{+}21]. Radio-based [WLL^{+}21]. Random [JTLL21]. Rate [GGBD20]. Reader [KHG^{+}20]. Real [DLT^{+}23]. Real-time [DLT^{+}23]. Realistic [GMT23]. Reality [ZUH^{+}22]. Recognition [LWL^{+}20, MAM^{+}21, WLS23, YXH^{+}22, YNS^{+}21, ZCC^{+}21, ZYJ^{+}23]. Reconfigurable [ADY^{+}21]. Reduce [LKR^{+}23]. Reinforcement [MAM^{+}21]. Relation [HH22]. Relationship [WLS23]. Relationships [AENM^{+}22]. Reliable [LXL^{+}23]. Representation [HAK22, HH22]. Requirements [ARP21]. Research [JPGM20]. Resilient [KKBL20]. Resource [ACS20, KWS22, MZC^{+}22, NG21, RVDA20, TFPNT23, XDP^{+}22]. Resource-Constrained [KWS22, ACS20]. Review [KK22, NPS^{+}22]. RF [SBR20]. RFID [ZYJ^{+}23]. RFID-single [ZYJ^{+}23]. Robust [AAN^{+}23, BK21, FAlS^{+}22, YXH^{+}22, ZJZ^{+}23]. RSSI [BA20]. RSSI-based [BA20]. Rubik [ARA23]. Runtime [LWLJ22, MZC^{+}22].

Safety [KSS21]. Sampling [GGBD20]. scale [LLL^{+}22]. Scenario [GMT23]. Scheme [AC20, SAGT20]. Science [BDJ^{+}22, DLR^{+}23]. Score [AAA^{+}21]. SCVS [MAD^{+}22]. SDN [BBS^{+}21, MCRP20]. SDN-Based [BBS^{+}21]. Seconds [MKC^{+}23]. Secure [GMHT22]. Security [ARP21, ARA23, EAM^{+}21, LKWM21]. Selection [KKY22]. Self [GTG^{+}21, KKR^{+}22]. Self-adaptive [KKR^{+}22]. Self-Updating [GTG^{+}21]. Semantic [GTG^{+}21, ND20]. Sensing [BK21, BMPP21, LMX^{+}20, MHL^{+}23, MKC^{+}23, PSLK21, Pic23, SAGT20, WLS23, ZCGL22]. Sensitive [XDP^{+}22]. Sensor [BJM21, HAK22, HST^{+}20, KKY22, PSW^{+}22, jWCS21]. Sensor-based [HST^{+}20]. Sensors [ATB^{+}22, BEM22]. Separability [BRL22]. Series [BRL22, GGBD20, HH22]. Service [HS21, LWL^{+}20]. Services [GGT^{+}23, MAD^{+}22]. Session [ARA23]. Shared [SBC^{+}22]. SiFi [GTG^{+}21]. Sign [ZJZ^{+}23]. Signature [LHJW23]. Simplifying [PDM^{+}20]. single [ZYJ^{+}23]. Smart [BEM22, CKCZY22, HGG^{+}23, JLZ^{+}20, KSBD20, LKWM21, LWL^{+}20, MAD^{+}22, PCMS21, SB21, SBC^{+}22, WWZZ20, ZCGL22, TPFF20]. Smart-city [MAD^{+}22]. Smartphone [ZZL^{+}23]. Smartphone-based [ZZL^{+}23]. Smartphones [jWCS21]. Snooping [ZZL^{+}23]. Soccer [HST^{+}20]. Social [JPGM20]. Sonar [ZZL^{+}23]. Source [BJM21, TF21]. sourced [ASD^{+}20]. Spaces [LKWM21, LWL^{+}20]. Spatial [BJM21]. Special [MHL^{+}23, Pic23]. Stack [KK22]. State [FAlS^{+}22, ZYJ^{+}23]. Stations [EAM^{+}21]. Stay [MKC^{+}23]. Storage [BMPP21]. Strategy [BRL22]. Street [LXML21]. Street-level [LXML21]. Study [AAA^{+}21, HGG^{+}23]. Sub [MKC^{+}23]. Sub-60 [MKC^{+}23]. subscribe [MGSD22]. Success [MBHF23]. Supervision [ZGX^{+}23]. supported [ASD^{+}20]. Surveillance [MAD^{+}22]. Survey [ARP21, DGL^{+}21, SDS^{+}22]. Swarm [AC20]. System [KQTC20, PSLK21, SBC^{+}22, TF21, YXH^{+}22, ZUH^{+}22]. Systematic [NPS^{+}22]. Systems [BBS^{+}21].

Tag [ZYJ^{+}23]. Tale [EAM^{+}21]. TargetFinder [KQTC20]. Targets [KQTC20]. Tensor [ZCC^{+}21]. Tensor-Compressed [ZCC^{+}21]. Testbed [JPGM20, TDD^{+}22]. Theory [DGL^{+}21]. Things [KKBL20, AWRP23, ARP21, AC20, AAA^{+}21, ADY^{+}21, BRL22, BJM21, CRCR21, GMHT22, HGG^{+}23, JTLL21, KKR^{+}22, KLMS22, MGSD22, NLQ22, ND20, TF21, ZUH^{+}22, DP20]. Tierless [LKR^{+}23]. Time [BRL22, GGBD20, HH22, MKC^{+}23, XDP^{+}22, DLT^{+}23]. Time-Sensitive [XDP^{+}22]. Timeliness [SBR20]. Timeout [PDM^{+}20]. Tiny [LLL^{+}22]. tools [DSS^{+}20]. Traces [jWCS21]. Tracing [TDD^{+}22]. Tracking [ACS20, DLT^{+}23]. Trading [ZGX^{+}23]. Traffic [KSS21]. Training [HST^{+}20]. Trajectory [WLS23]. Transactions [DP20, ZGX^{+}23]. Transformation [BDJ^{+}22, HAK22]. Transmission [BMPP21, KK21]. Transmissions [AEMK^{+}21]. Transport [KKY22]. Trials [EFO^{+}20]. TSCH [EFO^{+}20, GMT23]. Twin [PMZ23]. Two [EAM^{+}21].

UAVs [BMPM20, BK21]. Ubiquitous [BPR^{+}23]. UltraSnoop [ZZL^{+}23]. Ultrasonic [ZZL^{+}23]. Underwater [DLR^{+}23]. Unobtrusive [LWL^{+}20]. Unsupervised [KKY22]. Updating [GTG^{+}21]. Upscaling [DLR^{+}23]. User [LWL^{+}20, NG21, SLLC21, SBC^{+}22]. Using [ATB^{+}22, BK21, DLR^{+}23, JCY^{+}23, KKBL20, WEOR23, jWCS21, ZJZ^{+}23].

Value [KSBD20]. Vehicle [DLT^{+}23, EAM^{+}21]. Vehicles [KSS21]. Velocity [KSS21]. Verification [BEM22, LHJW23]. via [JLZ^{+}20, LWLJ22, LHJW23, LXL^{+}23, YXH^{+}22, ZZL^{+}23]. Vibration [MVC22]. Video [MAD^{+}22, ZCC^{+}21]. ViMo [ZJZ^{+}23]. VioLinn [TFPNT23]. Virtual [PF23]. Visibility [HNB^{+}20]. Visual [WLS23]. Vital [ZJZ^{+}23]. ViWise [WLS23].

Warning [AAA^{+}21]. WatchDog [DLT^{+}23]. Wave [LHJW23, LZW^{+}23]. Wearable [ATB^{+}22, BPR^{+}23, LMX^{+}20]. Web [KLMS22, ND20]. Who [SBC^{+}22]. WiFi [MAM^{+}21, SLLC21]. WiFi-Enabled [SLLC21]. Wild [EFO^{+}20]. Wireless [HNB^{+}20, LLL^{+}22, MHL^{+}23, Pic23, TDD^{+}22, TDB^{+}23, WLS23, YXH^{+}22, MCRP20]. WISE [AC20]. Word [Pic23]. Workflows [CKCZY22]. World [KHG^{+}20]. Worn [KHG^{+}20]. Wrist [KHG^{+}20]. Wrist-Worn [KHG^{+}20].

References

Anzanpour:2021:EAC

[AAA^{+}21] Arman Anzanpour, Delaram Amiri, Iman Azimi, Marco Levorato, Nikil Dutt, Pasi Liljeberg, and Amir M. Rahmani. Edge-assisted control for healthcare Internet of Things: a case study on PPG-based early warning score. ACM Transactions on Internet of Things (TIOT), 2(1):1:1–1:21, February 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                                   tronic). URL https://dl.acm.
                                                   org/doi/10.1145/3407091.

Ahmar:2023:DRM

[AAN^{+}23] Absar-Ul-Haque Ahmar, Emekcan Aras, Thien Duc Nguyen, Sam Michiels, Wouter Joosen, and Danny Hughes. Design of a robust MAC protocol for LoRa. ACM Transactions on Internet of Things (TIOT), 4(1): 3:1–3:??, February 2023. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3557048.

Ammar:2020:WLI

[AC20] Mahmoud Ammar and Bruno Crispo. WISE: a lightweight intelligent swarm attestation scheme for the Internet of Things. ACM Transactions on Internet of Things (TIOT), 1 (3):19:1–19:30, July 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3386688.

Aftab:2020:EOC

[ACS20] Muhammad Aftab, Sid Chi-Kin Chau, and Prashant Shenoy. Efficient online classification and tracking on resource-constrained IoT devices. ACM Transactions on Internet of Things (TIOT), 1 (3):20:1–20:29, July 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3392051.

Aras:2021:CLP

[ADY^{+}21] Emekcan Aras, Stéphane Delbruel, Fan Yang, Wouter Joosen, and Danny Hughes. Chimera: a low-power reconfigurable platform for Internet of Things. ACM Transactions on Internet of Things (TIOT), 2(2): 10:1–10:25, May 2021. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3440995.

AlNahas:2021:BCT

[AEMK^{+}21] Beshr Al Nahas, Antonio Escobar-Molero, Jirka Klaue, Simon Duquennoy, and Olaf Landsiedel. BlueFlood: Concurrent transmissions for multi-hop Bluetooth 5-modeling and evaluation. ACM Transactions on Internet of Things (TIOT), 2 (4):22:1–22:30, November 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3462755.

Apthorpe:2022:YMI

[AENM^{+}22] Noah Apthorpe, Pardis Emami-Naeini, Arunesh Mathur, Marshini Chetty, and Nick Feamster. You, me, and IoT: How Internetconnected consumer devices affect interpersonal relationships. ACM Transactions on Internet of Things (TIOT), 3(4):25:1– 25:??, November 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3539737.

Attkan:2023:RCC

[ARA23] Ankit Attkan, Virender Ranga, and Priyanka Ahlawat. A Rubik’s cube cryptosystem-based authentication and session key generation model driven in blockchain environment for IoT security. ACM Transactions on Internet of Things (TIOT), 4(2): 14:1–14:??, May 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586578.

Alhirabi:2021:SPR

[ARP21] Nada Alhirabi, Omer Rana, and Charith Perera. Security and privacy requirements for the Internet of Things: a survey. ACM Transactions on Internet of Things (TIOT), 2(1): 6:1–6:37, February 2021. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3437537.

Abdelaal:2020:MGS

[ASD^{+}20] Mohamed Abdelaal, Suriya Sekar, Frank Dürr, Kurt Rothermel, Susanne Becker, and Dieter Fritsch. MapSense: Grammarsupported inference of indoor objects from crowd-sourced 3D point clouds. ACM Transactions on Internet of Things (TIOT), 1 (2):11:1–11:28, April 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3379342.

An:2022:MMB

[ATB^{+}22] Sizhe An, Yigit Tuncel, Toygun Basaklar, Gokul K. Krishnakumar, Ganapati Bhat, and Umit Y. Ogras. MGait: Modelbased gait analysis using wearable bend and inertial sensors. ACM Transactions on Internet of Things (TIOT), 3(1): 7:1–7:24, February 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3485434.

AlMuhander:2023:IPM

[AWRP23] Bayan Al Muhander, Jason Wiese, Omer Rana, and Charith Perera. Interactive privacy management: Toward enhancing privacy awareness and control in the Internet of Things. ACM Transactions on Internet of Things (TIOT), 4(3):18:1– 18:??, August 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3600096.

Bauer:2020:TLC

[BA20] Jan Bauer and Nils Aschenbruck. Towards a low-cost RSSI-based crop monitoring. ACM Transactions on Internet of Things (TIOT), 1(4):21:1– 21:26, October 2020. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3393667.

Bouloukakis:2021:PDE

[BBS^{+}21] Georgios Bouloukakis, Kyle Benson, Luca Scalzotto, Paolo Bellavista, Casey Grant, Valérie Issarny, Sharad Mehrotra, Ioannis Moscholios, and Nalini Venkatasubramanian. PrioDeX: a data exchange middleware for efficient event prioritization in SDN-based IoT systems. ACM Transactions on Internet of Things (TIOT), 2(3): 19:1–19:32, July 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3456301.

Brauner:2022:CSP

[BDJ^{+}22] Philipp Brauner, Manuela Dalibor, Matthias Jarke, Ike Kunze, István Koren, Gerhard Lakemeyer, Martin Liebenberg, Judith Michael, Jan Pennekamp, Christoph Quix, Bernhard Rumpe, Wil van der Aalst, Klaus Wehrle, Andreas Wortmann, and Martina Ziefle. A computer science perspective on digital transformation in production. ACM Transactions on Internet of Things (TIOT), 3(2): 15:1–15:32, May 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3502265.

Birnbach:2022:HHP

[BEM22] Simon Birnbach, Simon Eberz, and Ivan Martinovic. Haunted house: Physical smart home event verification in the presence of compromised sensors. ACM Transactions on Internet of Things (TIOT), 3(3): 18:1–18:28, August 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3506859.

Bradbury:2021:SSL

[BJM21] Matthew Bradbury, Arshad Jhumka, and Carsten Maple. A spatial source location privacyaware duty cycle for Internet of Things sensor networks. ACM Transactions on Internet of Things (TIOT), 2(1):4:1– 4:32, February 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3430379.

Boubrima:2021:RES

[BK21] Ahmed Boubrima and Edward W. Knightly. Robust environmental sensing using UAVs. ACM Transactions on Internet of Things (TIOT), 2(4):25:1– 25:20, November 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3464943.

Bhattacharya:2020:IIB

[BMPM20] Debarpan Bhattacharya, Sudip Misra, Nidhi Pathak, and Anandarup Mukherjee. IDeA: IoTbased autonomous aerial demarcation and path planning for precision agriculture with UAVs. ACM Transactions on Internet of Things (TIOT), 1 (3):16:1–16:21, July 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3379930.

Buddhika:2021:LED

[BMPP21] Thilina Buddhika, Matthew Malensek, Shrideep Pallickara, and Sangmi Lee Pallickara. Living on the edge: Data transmission, storage, and analytics in continuous sensing environments. ACM Transactions on Internet of Things (TIOT), 2(3): 16:1–16:31, July 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3450767.

Bannis:2023:IMF

[BPR^{+}23] Adeola Bannis, Shijia Pan, Carlos Ruiz, John Shen, Hae Young Noh, and Pei Zhang. IDIoT: Multimodal framework for ubiquitous identification and assignment of human-carried wearable devices. ACM Transactions on Internet of Things (TIOT), 4(2):11:1–11:??, May 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579832.

Borges:2022:CSI

[BRL22] João B. Borges, Heitor S.

Ramos, and Antonio A. F. Loureiro. A classification strategy for Internet of Things data based on the class separability analysis of time series dynamics. ACM Transactions on Internet of Things (TIOT), 3(3):23:1– 23:30, August 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3533049.

Constantinou:2022:GPI

[CKCZY22] Soteris Constantinou, Andreas Konstantinidis, Panos K. Chrysanthis, and Demetrios Zeinalipour-Yazti. Green planning of IoT home automation workflows in smart buildings. ACM Transactions on Internet of Things (TIOT), 3(4):29:1–29:??, November 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3549549.

Chowdhury:2021:NIA

[CRCR21] Morshed Chowdhury, Biplob Ray, Sujan Chowdhury, and Sutharshan Rajasegarar. A novel insider attack and machine learning based detection for the Internet of Things. ACM Transactions on Internet of Things (TIOT), 2(4):26:1– 26:23, November 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3466721.

Dhar:2021:SDM

[DGL^{+}21] Sauptik Dhar, Junyao Guo, Jiayi (Jason) Liu, Samarth Tripathi, Unmesh Kurup, and Mohak Shah. A survey of ondevice machine learning: an algorithms and learning theory perspective. ACM Transactions on Internet of Things (TIOT), 2(3):15:1–15:49, July 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3450494.

Dar:2023:UFC

[DLR^{+}23] Farooq Dar, Mohan Liyanage, Marko Radeta, Zhigang Yin, Agustin Zuniga, Sokol Kosta, Sasu Tarkoma, Petteri Nurmi, and Huber Flores. Upscaling fog computing in oceans for underwater pervasive data science using low-cost microclouds. ACM Transactions on Internet of Things (TIOT), 4 (2):9:1–9:??, May 2023. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3575801.

Dong:2023:WRT

[DLT^{+}23] Zheng Dong, Yan Lu, Guangmo Tong, Yuanchao Shu, Shuai Wang, and Weisong Shi. Watch-Dog: Real-time vehicle tracking on geo-distributed edge nodes. ACM Transactions on Internet of Things (TIOT), 4(1): 2:1–2:??, February 2023. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549551.

Dustdar:2020:ATI

[DP20] Schahram Dustdar and Gian Pietro Picco. ACM Transactions on Internet of Things: Inaugural issue editorial. ACM Transactions on Internet of Things (TIOT), 1 (1):1:1–1:2, February 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3379599.

Prado:2020:BAP

[DSS^{+}20] Miguel De Prado, Jing Su, Rabia Saeed, Lorenzo Keller, Noelia Vallez, Andrew Anderson, David Gregg, Luca Benini, Tim Llewellynn, Nabil Ouerhani, Rozenn Dahyot, and Nuria Pazos. Bonseyes AI pipelinebringing AI to you: End-to-end integration of data, algorithms, and deployment tools. ACM Transactions on Internet of Things (TIOT), 1(4):26:1– 26:25, October 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3403572.

ElHussini:2021:TTE

[EAM^{+}21] Hossam ElHussini, Chadi Assi, Bassam Moussa, Ribal Atallah, and Ali Ghrayeb. A tale of two entities: Contextualizing the security of electric vehicle charging stations on the power grid. ACM Transactions on Internet of Things (TIOT), 2(2):8:1–8:21, May 2021. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3437258.

Elsts:2020:TNH

[EFO^{+}20] Atis Elsts, Xenofon Fafoutis, George Oikonomou, Robert Piechocki, and Ian Craddock. TSCH networks for health IoT: Design, evaluation, and trials in the wild. ACM Transactions on Internet of Things (TIOT), 1 (2):9:1–9:27, April 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3366617.

Fomichev:2022:NRC

[FAlS^{+}22] Mikhail Fomichev, Luis F.

Abanto-leon, Max Stiegler, Alejandro Molina, Jakob Link, and Matthias Hollick. Next2You: Robust copresence detection based on channel state information. ACM Transactions on Internet of Things (TIOT), 3(2): 11:1–11:31, May 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3491244.

Gupta:2020:DCB

[GGBD20] Ashish Gupta, Hari Prabhat Gupta, Bhaskar Biswas, and Tanima Dutta. A divide-and-conquer-based early classification approach for multivariate time series with different sampling rate components in IoT. ACM Transactions on Internet of Things (TIOT), 1(2): 10:1–10:21, April 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3375877.

Gu:2023:HCE

[GGT^{+}23] Siyuan Gu, Deke Guo, Guoming Tang, Lailong Luo, Yuchen Sun, and Xueshan Luo. HyEdge: a cooperative edge computing framework for provisioning private and public services. ACM Transactions on Internet of Things (TIOT), 4(2):

13:1–13:??, May 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3585078.

Gunnarsson:2022:PEG

[GMHT22] Martin Gunnarsson, Krzysztof Mateusz Malarski, Rikard Höglund, and Marco Tiloca. Performance evaluation of group OSCORE for secure group communication in the Internet of Things. ACM Transactions on Internet of Things (TIOT), 3(3):19:1– 19:31, August 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3523064.

Gonzalez:2023:CPC

[GMT23] Ivonne Andrea Mantilla Gonzalez, Florian Meyer, and Volker Turau. A comprehensive performance comparison of IEEE 802.15.4 DSME and TSCH in a realistic IoT scenario for industrial applications. ACM Transactions on Internet of Things (TIOT), 4(3):16:1–16:??, August 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3595188.

Guo:2021:SSU

[GTG^{+}21] Deke Guo, Xiaoqiang Teng, Yulan Guo, Xiaolei Zhou, and Zhong Liu. SiFi: Self-updating of indoor semantic floorplans for annotated objects. ACM Transactions on Internet of Things (TIOT), 2(3):17:1–17:21, July 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3450567.

Hajihassani:2022:ASD

[HAK22] Omid Hajihassani, Omid Ardakanian, and Hamzeh Khazaei. Anonymizing sensor data on the edge: a representation learning and transformation approach. ACM Transactions on Internet of Things (TIOT), 3(1):8:1–8:26, February 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485820.

Hamed:2023:QIS

[HGG^{+}23] Naeima Hamed, Andrea Gaglione, Alex Gluhak, Omer Rana, and Charith Perera. Query interface for smart city Internet of Things data marketplaces: a case study. ACM Transactions on Internet of Things (TIOT), 4(3):19:1– 19:??, August 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3609336.

Hussein:2022:DAR

[HH22] Amir Hussein and Hazem Hajj.

Domain adaptation with representation learning and nonlinear relation for time series. ACM Transactions on Internet of Things (TIOT), 3(2): 12:1–12:26, May 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3502905.

Huang:2020:CWC

[HNB^{+}20] Hua Huang, Chien-Chun Ni, Xiaomeng Ban, Andrew T. Schneider, Jie Gao, and Shan Lin. Connected wireless camera network deployment with visibility coverage. ACM Transactions on Internet of Things (TIOT), 1(4):25:1–25:19, October 2020. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3399431.

Huang:2021:ESC

[HS21] Chih-Kai Huang and Shan-Hsiang Shen. Enabling service cache in edge clouds. ACM Transactions on Internet of Things (TIOT), 2(3): 18:1–18:24, July 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3456564.

Haladjian:2020:SBD

[HST^{+}20] Juan Haladjian, Daniel Schlabbers, Sajjad Taheri, Max Tharr, and Bernd Bruegge. Sensorbased detection and classification of soccer goalkeeper training exercises. ACM Transactions on Internet of Things (TIOT), 1 (2):12:1–12:20, April 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3372342.

Juncen:2023:MFG

[JCY^{+}23] Zhu Juncen, Jiannong Cao, Yanni Yang, Wei Ren, and Huizi Han. mmDrive: Finegrained fatigue driving detection using mmWave radar. ACM Transactions on Internet of Things (TIOT), 4(4):26:1– 26:??, November 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3614437.

Ji:2020:ASH

[JLZ^{+}20] Xiaoyu Ji, Chaohao Li, Xinyan Zhou, Juchuan Zhang, Yanmiao Zhang, and Wenyuan Xu. Authenticating smart home devices via home limited channels. ACM Transactions on Internet of Things (TIOT), 1 (4):24:1–24:24, October 2020. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3399432.

Jiang:2020:MIT

[JPGM20] Jie Jiang, Riccardo Pozza, Nigel Gilbert, and Klaus Moessner. MakeSense: an IoT testbed for social research of indoor activities. ACM Transactions on Internet of Things (TIOT), 1 (3):17:1–17:25, July 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3381914.

Jiang:2021:LPP

[JTLL21] Linshan Jiang, Rui Tan, Xin Lou, and Guosheng Lin. On lightweight privacy-preserving collaborative learning for Internet of Things by independent random projections. ACM Transactions on Internet of Things (TIOT), 2(2):11:1– 11:32, May 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3441303.

Wu:2021:CQC

[jWCS21] Fang jing Wu, Ying-Jun Chen, and Sok-Ian Sou. CoCo: Quantifying correlations between mobility traces using sensor data from smartphones. ACM Transactions on Internet of Things (TIOT), 2(3):20:1–20:22, July 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3457139.

Kiaghadi:2020:CMI

[KHG^{+}20] Ali Kiaghadi, Pan Hu, Jeremy Gummeson, Soha Rostaminia, and Deepak Ganesan. Continuous measurement of interactions with the physical world with a wrist-worn backscatter reader. ACM Transactions on Internet of Things (TIOT), 1 (2):7:1–7:22, April 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3375800.

Kalita:2021:OTC

[KK21] Alakesh Kalita and Manas Khatua. Opportunistic transmission of control packets for faster formation of 6TiSCH network. ACM Transactions on Internet of Things (TIOT), 2(1):5:1–5:29, February 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3430380.

Kalita:2022:IEO

[KK22] Alakesh Kalita and Manas Khatua. 6TiSCH — IPv6 enabled open stack IoT network formation: a review. ACM Transactions on Internet of Things (TIOT), 3(3):24:1– 24:36, August 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3536166.

Kim:2020:RAA

[KKBL20] Hokeun Kim, Eunsuk Kang, David Broman, and Edward A. Lee. Resilient authentication and authorization for the Internet of Things (IoT) using edge computing. ACM Transactions on Internet of Things (TIOT), 1(1):4:1–4:27, February 2020. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3375837.

Kirchhof:2022:MDS

[KKR^{+}22] Jörg Christian Kirchhof, Anno Kleiss, Bernhard Rumpe, David Schmalzing, Philipp Schneider, and Andreas Wortmann. Model-driven self-adaptive deployment of Internet of Things applications with automated modification proposals. ACM Transactions on Internet of Things (TIOT), 3(4):30:1–30:??, November 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3549553.

Killeen:2022:UDS

[KKY22] Patrick Killeen, Iluju Kiringa, and Tet Yeap. Unsupervised dynamic sensor selection for IoT-Based predictive maintenance of a fleet of public transport buses. ACM Transactions on Internet of Things (TIOT), 3(3): 21:1–21:36, August 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3530991.

Krishna:2022:DDE

[KLMS22] Ajay Krishna, Michel Le Pallec, Radu Mateescu, and Gwen Salaün. Design and deployment of expressive and correct Web of Things applications. ACM Transactions on Internet of Things (TIOT), 3(1): 1:1–1:30, February 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3475964.

Khazbak:2020:TPP

[KQTC20] Youssef Khazbak, Junpeng Qiu, Tianxiang Tan, and Guohong Cao. TargetFinder: a privacy preserving system for locating targets through IoT cameras. ACM Transactions on Internet of Things (TIOT), 1 (3):14:1–14:23, July 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3375878.

Khamesi:2020:PVD

[KSBD20] Atieh R. Khamesi, Simone Silvestri, D. A. Baker, and Alessandra De Paola. Perceived-value-driven optimization of energy consumption in smart homes. ACM Transactions on Internet of Things (TIOT), 1(2): 13:1–13:26, April 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3375801.

Kang:2021:VOP

[KSS21] Liuwang Kang, Ankur Sarker, and Haiying Shen. Velocity optimization of pure electric vehicles with traffic dynamics and driving safety considerations. ACM Transactions on Internet of Things (TIOT), 2(1): 7:1–7:24, February 2021. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3433678.

Kumar:2022:NAC

[KWS22] Arun Kumar, Zhijie Wang, and Abhishek Srivastava. A novel approach for classification in resource-constrained environments. ACM Transactions on Internet of Things (TIOT), 3 (4):32:1–32:??, November 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3549552.

Li:2023:MAH

[LHJW23] Wanqing Li, Tongtong He, Nan Jing, and Lin Wang. mmHSV: In-air handwritten signature verification via millimeter-wave radar. ACM Transactions on Internet of Things (TIOT), 4 (4):27:1–27:??, November 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3614443.

Lubbers:2023:CTL

[LKR^{+}23] Mart Lubbers, Pieter Koopman, Adrian Ramsingh, Jeremy Singer, and Phil Trinder. Could tierless languages reduce IoT development grief? ACM Transactions on Internet of Things (TIOT), 4(1):6:1–6:??, February 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3572901.

Lin:2021:ASM

[LKWM21] Changyuan Lin, Hamzeh Khazaei, Andrew Walenstein, and Andrew Malton. Autonomic security management for IoT smart spaces. ACM Transactions on Internet of Things (TIOT), 2(4):27:1–27:20, November 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3466696.

Lu:2022:MSF

[LLL^{+}22] Feng Lu, Wei Li, Song Lin, Chengwangli Peng, Zhiyong Wang, Bin Qian, Rajiv Ranjan, Hai Jin, and Albert Y. Zomaya. Multi-scale features fusion for the detection of tiny bleeding in wireless capsule endoscopy images. ACM Transactions on Internet of Things (TIOT), 3(1):2:1–2:19, February 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477540.

Lan:2020:CBA

[LMX^{+}20] Guohao Lan, Dong Ma, Weitao Xu, Mahbub Hassan, and Wen Hu. Capacitor-based activity sensing for kineticpowered wearable IoTs. ACM Transactions on Internet of Things (TIOT), 1(1):2:1–2:26, February 2020. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3362124.

Luo:2020:GRS

[LWL^{+}20] Chengwen Luo, Jiawei Wu, Jianqiang Li, Jia Wang, Weitao Xu, Zhong Ming, Bo Wei, Wei Li, and Albert Y. Zomaya. Gait recognition as a service for unobtrusive user identification in smart spaces. ACM Transactions on Internet of Things (TIOT), 1(1): 5:1–5:21, February 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3375799.

Li:2022:MRM

[LWLJ22] Yifu Li, Lening Wang, Dongyoon Lee, and Ran Jin. Monitoring runtime metrics of fog manufacturing via a qualitative and quantitative (QQ) control chart. ACM Transactions on Internet of Things (TIOT), 3 (2):14:1–14:19, May 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3501262.

Luo:2023:CMA

[LXL^{+}23] Junzhou Luo, Zhuqing Xu, Jingkai Lin, Ciyuan Chen, and Runqun Xiong. CH-MAC: Achieving low-latency reliable communication via coding and hopping in LPWAN. ACM Transactions on Internet of Things (TIOT), 4(4):24:1– 24:??, November 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3617505.

Li:2021:LSL

[LXML21] Ruixiang Li, Rui Xu, Yuanyuan Ma, and Xiangyang Luo. LandmarkMiner: Street-level network landmarks mining method for IP geolocation. ACM Transactions on Internet of Things (TIOT), 2(3):21:1–21:22, July 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3457409.

Liang:2023:AMY

[LZW^{+}23] Yumeng Liang, Anfu Zhou, Xinzhe Wen, Wei Huang, Pu Shi, Lingyu Pu, Huanhuan Zhang, and Huadong Ma. airBP: Monitor your blood pressure with millimeter-wave in the air. ACM Transactions on Internet of Things (TIOT), 4(4):28:1– 28:??, November 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3614439.

Myneni:2022:SAE

[MAD^{+}22] Sowmya Myneni, Garima Agrawal, Yuli Deng, Ankur Chowdhary, Neha Vadnere, and Dijiang Huang. SCVS: On AI and edge clouds enabled privacy-preserved smart-city video surveillance services. ACM Transactions on Internet of Things (TIOT), 3 (4):28:1–28:??, November 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3542953.

Ma:2021:LPI

[MAM^{+}21] Yongsen Ma, Sheheryar Arshad, Swetha Muniraju, Eric Torkildson, Enrico Rantala, Klaus Doppler, and Gang Zhou. Location- and personindependent activity recognition with WiFi, deep neural networks, and reinforcement learning. ACM Transactions on Internet of Things (TIOT), 2(1):3:1– 3:25, February 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3424739.

Moedt:2023:EIP

[MBHF23] Wouter Moedt, Reinhard Bernsteiner, Margeret Hall, and Ann Fruhling. Enhancing IoT project success through agile best practices. ACM Transactions on Internet of Things (TIOT), 4(1): 5:1–5:??, February 2023. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3568170.

Murillo:2020:AWS

[MCRP20] Yuri Murillo, Alessandro Chiumento, Brecht Reynders, and Sofie Pollin. An all-wireless SDN framework for BLE mesh. ACM Transactions on Internet of Things (TIOT), 1(4):27:1– 27:30, October 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3403581.

Montori:2022:MLA

[MGSD22] Federico Montori, Lorenzo Gigli, Luca Sciullo, and Marco Di Felice. LA-MQTT: Locationaware publish-subscribe communications for the Internet of Things. ACM Transactions on Internet of Things (TIOT), 3(3):20:1–20:28, August 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3529978.

Ma:2023:ISI

[MHL^{+}23] Huadong Ma, Yuan He, Mo Li, Neal Patwari, and Stephan Sigg. Introduction to the special issue on wireless sensing for IoT. ACM Transactions on Internet of Things (TIOT), 4(4):21:1– 21:??, November 2023. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3633078.

Mandal:2023:EMM

[MKC^{+}23] Ratna Mandal, Prasenjit Karmakar, Soumyajit Chatterjee, Debaleen Das Spandan, Shouvit Pradhan, Sujoy Saha, Sandip Chakraborty, and Subrata Nandi. Exploiting multimodal contextual sensing for city-bus’s stay location characterization: Towards sub-60 seconds accurate arrival time prediction. ACM Transactions on Internet of Things (TIOT), 4(1):1:1–1:??, February 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549548.

Michala:2022:VEC

[MVC22] Anna Lito Michala, Ioannis Vourganas, and Andrea Coraddu. Vibration edge computing in maritime IoT. ACM Transactions on Internet of Things (TIOT), 3(1):6:1– 6:18, February 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3484717.

Miele:2022:RRM

[MZC^{+}22] Antonio Miele, Henry Zárate, Luca Cassano, Cristiana Bolchini, and Jorge E. Ortiz. A runtime resource management and provisioning middleware for fog computing infrastructures. ACM Transactions on Internet of Things (TIOT), 3(3): 17:1–17:29, August 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3506718.

Norris:2022:IFF

[NCV^{+}22] Michael Norris, Z. Berkay Celik, Prasanna Venkatesh, Shulin Zhao, Patrick McDaniel, Anand Sivasubramaniam, and Gang Tan. IoTRepair: Flexible fault handling in diverse IoT deployments. ACM Transactions on Internet of Things (TIOT), 3(3):22:1–22:33, August 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3532194.

Novo:2020:SII

[ND20] Oscar Novo and Mario Di Francesco. Semantic interoperability in the IoT: Extending the Web of Things architecture. ACM Transactions on Internet of Things (TIOT), 1(1): 6:1–6:25, February 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3375838.

Naha:2021:MCB

[NG21] Ranesh Kumar Naha and Saurabh Garg. Multi-criteria-based dynamic user behaviouraware resource allocation in fog computing. ACM Transactions on Internet of Things (TIOT), 2(1):2:1–2:31, February 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3423332.

Ngo:2022:AAD

[NLQ22] Mao V. Ngo, Tie Luo, and Tony Q. S. Quek. Adaptive anomaly detection for Internet of Things in hierarchical edge computing: a contextualbandit approach. ACM Transactions on Internet of Things (TIOT), 3(1):4:1–4:23, February 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480172.

Neha:2022:SRO

[NPS^{+}22] Benazir Neha, Sanjaya Kumar Panda, Pradip Kumar Sahu, Kshira Sagar Sahoo, and Amir H. Gandomi. A systematic review on osmotic computing. ACM Transactions on Internet of Things (TIOT), 3(2): 9:1–9:30, May 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3488247.

Piva:2021:EDC

[PCMS21] Mauro Piva, Andrea Coletta, Gaia Maselli, and John A. Stankovic. Environment-driven communication in battery-free smart buildings. ACM Transactions on Internet of Things (TIOT), 2(2):14:1–14:30, May 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3448739.

Peros:2020:SCA

[PDM^{+}20] Stefanos Peros, Sté;phane Delbruel, Sam Michiels, Wouter Joosen, and Danny Hughes. Simplifying CPS application development through fine-grained, automatic timeout predictions. ACM Transactions on Internet of Things (TIOT), 1(3): 18:1–18:30, July 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3385960.

Picano:2023:CAF

[PF23] Benedetta Picano and Romano Fantacci. A channel-aware FL approach for virtual machine placement in 6G edge intelligent ecosystems. ACM Transactions on Internet of Things (TIOT), 4(2):12:1–12:??, May 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3584705.

Picco:2023:SIW

[Pic23] Gian Pietro Picco. Special issue on wireless sensing for IoT: a word from the Editorin-Chief. ACM Transactions on Internet of Things (TIOT), 4 (4):20:1–20:??, November 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3633752.

Picone:2023:FMA

[PMZ23] Marco Picone, Marco Mamei, and Franco Zambonelli. A flexible and modular architecture for edge digital twin: Implementation and evaluation. ACM Transactions on Internet of Things (TIOT), 4(1):8:1– 8:??, February 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3573206.

Petrolo:2021:ASG

[PSLK21] Riccardo Petrolo, Zhambyl Shaikhanov, Yingyan Lin, and Edward Knightly. ASTRO: a system for off-grid networked drone sensing missions. ACM Transactions on Internet of Things (TIOT), 2(4):24:1–24:22, November 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3464942.

Panwar:2022:INA

[PSW^{+}22] Nisha Panwar, Shantanu Sharma, Guoxi Wang, Sharad Mehrotra, Nalini Venkatasubramanian, Mamadou H. Diallo, and Ardalan Amiri Sani. IoT notary: Attestable sensor data capture in IoT environments. ACM Transactions on Internet of Things (TIOT), 3(1):3:1–3:30, February 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3478290.

Righetti:2020:EDR

[RVDA20] Francesca Righetti, Carlo Vallati, Sajal K. Das, and Giuseppe Anastasi. An evaluation of the 6TiSCH distributed resource management mode. ACM Transactions on Internet of Things (TIOT), 1(4):23:1–23:31, October 2020. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3395927.

Sigrist:2020:HAO

[SAGT20] Lukas Sigrist, Rehan Ahmed, Andres Gomez, and Lothar Thiele. Harvesting-aware optimal communication scheme for infrastructure-less sensing. ACM Transactions on Internet of Things (TIOT), 1(4):22:1– 22:26, October 2020. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3395928.

Sater:2021:FLA

[SB21] Raed Abdel Sater and A. Ben Hamza. A federated learning approach to anomaly detection in smart buildings. ACM Transactions on Internet of Things (TIOT), 2(4):28:1– 28:23, November 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3467981.

Sikder:2022:WCM

[SBC^{+}22] Amit Kumar Sikder, Leonardo Babun, Z. Berkay Celik, Hidayet Aksu, Patrick McDaniel, Engin Kirda, and A. Selcuk Uluagac. Who’s controlling my device? Multi-user multidevice-aware access control system for shared smart home environment. ACM Transactions on Internet of Things (TIOT), 3(4):27:1–27:??, November 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543513.

Spork:2020:ITB

[SBR20] Michael Spörk, Carlo Alberto Boano, and Kay Römer. Improving the timeliness of Bluetooth low energy in dynamic RF environments. ACM Transactions on Internet of Things (TIOT), 1(2):8:1–8:32, April 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3375836.

Safi:2022:SIP

[SDS^{+}22] Miraqa Safi, Sajjad Dadkhah, Farzaneh Shoeleh, Hassan Mahdikhani, Heather Molyneaux, and Ali A. Ghorbani. A survey on IoT profiling, fingerprinting, and identification. ACM Transactions on Internet of Things (TIOT), 3(4):26:1– 26:??, November 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3539736.

Shi:2021:WEU

[SLLC21] Cong Shi, Jian Liu, Hongbo Liu, and Yingying Chen. WiFienabled user authentication through deep learning in daily activities. ACM Transactions on Internet of Things (TIOT), 2(2): 13:1–13:25, May 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3448738.

Trub:2023:LEE

[TDB^{+}23] Roman Trüb, Reto Da Forno, Andreas Biri, Jan Beutel, and Lothar Thiele. LSR: Energyefficient multi-modulation communication for inhomogeneous wireless IoT networks. ACM Transactions on Internet of Things (TIOT), 4(2):10:1–10:??, May 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579366.

Trub:2022:NID

[TDD^{+}22] Roman Trüb, Reto Da Forno, Lukas Daschinger, Andreas Biri, Jan Beutel, and Lothar Thiele. Non-intrusive distributed tracing of wireless IoT devices with the FlockLab 2 testbed. ACM Transactions on Internet of Things (TIOT), 3(1):5:1– 5:31, February 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3480248.

Turchet:2021:EAO

[TF21] Luca Turchet and Carlo Fischione. Elk audio OS: an open source operating system for the Internet of Musical Things. ACM Transactions on Internet of Things (TIOT), 2(2): 12:1–12:18, May 2021. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3446393.

Tocze:2023:VPA

[TFPNT23] Klervie Toczé, Ali J. Fahs, Guillaume Pierre, and Simin Nadjm-Tehrani. VioLinn: Proximityaware edge placementwith dynamic and elastic resource provisioning. ACM Transactions on Internet of Things (TIOT), 4(1): 7:1–7:??, February 2023. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3573125.

Turchet:2020:CSM

[TPFF20] Luca Turchet, Johan Pauwels, Carlo Fischione, and György Fazekas. Cloud-smart musical instrument interactions: Querying a large music collection with a smart guitar. ACM Transactions on Internet of Things (TIOT), 1(3): 15:1–15:29, July 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3377881.

Wang:2023:FID

[WEOR23] Han Wang, David Eklund, Alina Oprea, and Shahid Raza. FL4IoT: IoT device fingerprinting and identification using federated learning. ACM Transactions on Internet of Things (TIOT), 4(3):17:1–17:??, August 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3603257.

Wei:2021:NND

[WLL^{+}21] Bo Wei, Kai Li, Chengwen Luo, Weitao Xu, Jin Zhang, and Kuan Zhang. No need of data pre-processing: a general framework for radio-based device-free context awareness. ACM Transactions on Internet of Things (TIOT), 2(4):29:1– 29:26, November 2021. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3467980.

Wu:2023:VFV

[WLS23] Fang-Jing Wu, Sheng-Wun Lai, and Sok-Ian Sou. ViWise: Fusing visual and wireless sensing data for trajectory relationship recognition. ACM Transactions on Internet of Things (TIOT), 4(4):23:1–23:??, November 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3614441.

Wang:2020:PLS

[WWZZ20] Weizheng Wang, Qing Wang, Junwei Zhang, and Marco Zuniga. PassiveVLP: Leveraging smart lights for passive positioning. ACM Transactions on Internet of Things (TIOT), 1(1): 3:1–3:24, February 2020. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3362123.

Xavier:2022:MHT

[XDP^{+}22] Tiago C. S. Xavier, Flavia C.

Delicato, Paulo F. Pires, Claudio L. Amorim, Wei Li, and Albert Zomaya. Managing heterogeneous and time-sensitive IoT applications through collaborative and energy-aware resource allocation. ACM Transactions on Internet of Things (TIOT), 3 (2):10:1–10:28, May 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3488248.

Ye:2021:CAR

[YNS^{+}21] Juan Ye, Pakawat Nakwijit, Martin Schiemer, Saurav Jha, and Franco Zambonelli. Continual activity recognition with generative adversarial networks. ACM Transactions on Internet of Things (TIOT), 2(2): 9:1–9:25, May 2021. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3440036.

Yang:2022:QRM

[YXH^{+}22] Deliang Yang, Guoliang Xing, Jun Huang, Xiangmao Chang, and Xiaofan Jiang. QID: Robust mobile device recognition via a multi-coil Qi-wireless charging system. ACM Transactions on Internet of Things (TIOT), 3 (2):13:1–13:27, May 2022. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3498904.

Zhen:2021:FVF

[ZCC^{+}21] Peining Zhen, Hai-Bao Chen, Yuan Cheng, Zhigang Ji, Bin Liu, and Hao Yu. Fast video facial expression recognition by a deeply tensor-compressed LSTM neural network for mobile devices. ACM Transactions on Internet of Things (TIOT), 2 (4):23:1–23:26, November 2021. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3464941.

Zhu:2022:IIS

[ZCGL22] Hanwei Zhu, Sid Chi-Kin Chau, Gladhi Guarddin, and Weifa Liang. Integrating IoT-sensing and crowdsensing with privacy: Privacy-preserving hybrid sensing for smart cities. ACM Transactions on Internet of Things (TIOT), 3(4):31:1– 31:??, November 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

https://dl.acm.org/doi/10.
1145/3549550.

Zhao:2023:CLH

[ZGX^{+}23] Liushun Zhao, Deke Guo, Junjie Xie, Lailong Luo, and Yulong Shen. A closed-loop hybrid supervision framework of cryptocurrency transactions for data trading in IoT. ACM Transactions on Internet of Things (TIOT), 4(1):4:1–4:??, February 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3568171.

Zhang:2023:PVP

[ZJZ^{+}23] Bo Zhang, Boyu Jiang, Rong Zheng, Xiaoping Zhang, Jun Li, and Qiang Xu. Pi-ViMo: Physiology-inspired robust vital sign monitoring using mmWave radars. ACM Transactions on Internet of Things (TIOT), 4 (2):15:1–15:??, May 2023. CO-DEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3589347.

Zhang:2022:MAR

[ZUH^{+}22] Huanle Zhang, Mostafa Uddin, Fang Hao, Sarit Mukherjee, and Prasant Mohapatra. MAIDE: Augmented reality (AR)-facilitated mobile system for onboarding of Internet of Things (IoT) devices at ease. ACM Transactions on Internet of Things (TIOT), 3(2):16:1– 16:21, May 2022. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3506667.

Zhu:2023:MMO

[ZYJ^{+}23] Biaokai Zhu, Zejiao Yang, Yupeng Jia, Shengxin Chen, Jie Song, Sanman Liu, Ping Li, Feng Li, and Deng-Ao Li. MFD: Multi-object frequency feature recognition and state detection based on RFID-single tag. ACM Transactions on Internet of Things (TIOT), 4(4):25:1– 25:??, November 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3615665.

Zhao:2023:UPA

[ZZL^{+}23] Yanchao Zhao, Yiming Zhao, Si Li, Hao Han, and Lei Xie. UltraSnoop: Placementagnostic keystroke snooping via smartphone-based ultrasonic sonar. ACM Transactions on Internet of Things (TIOT), 4 (4):22:1–22:??, November 2023. CODEN ???? ISSN 2691-1914 (print), 2577-6207 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3614440.